Devante flies a plane against a headwind for 3168 miles. The return trip with the wind took 18 hours less time. If the wind speed is 9 mph, how fast does Devante fly the plane when there is no wind?
Answer by TimothyLamb(4379)

s = d / t
d = s * t
---
outbound:
3168 = (s - w)T
---
return:
3168 = (s + w)(T - 18)
---
(s - w)T = (s + w)(T - 18)
(s - 9)T = (s + 9)(T - 18)
sT - 9T = sT - 18s + 9T - 162
0 = -18s + 18T - 162
18T - 18s = 162
18(T - s) = 162
T - s = 9
s = T - 9
---
3168 = (s - w)T
3168 = (s - 9)T
3168 = (T - 9 - 9)T
3168 = (T - 18)T
TT - 18T - 3168 = 0
T^2 - 18T - 3168 = 0
---
the above quadratic equation is in standard form, with a=1, b=-18, and c=-3168
---
to solve the quadratic equation, by using the quadratic formula, plug this:
1 -18 -3168
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the two real roots (i.e. the two x-intercepts), of the quadratic are:
T = 66
T = -48
---
the negative root doesn't make sense for time, so use the positive root:
---
T = 66 hours
s = T - 9
s = 66 - 9
---
answer:
airplane speed in calm air: s = 57 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php